# Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers

### v1

@@ Abstract
In the era of generative AI, deep generative models (DGMs) with latent representations have gained tremendous popularity. Despite their impressive empirical performance, the statistical properties of these models remain underexplored. DGMs are often overparametrized, non-identifiable, and uninterpretable black boxes, raising serious concerns when deploying them in high-stakes applications. Motivated by this, we propose an interpretable deep generative modeling framework for rich data types with discrete latent layers, called Deep Discrete Encoders (DDEs). A DDE is a directed graphical model with multiple binary latent layers. Theoretically, we propose transparent identifiability conditions for DDEs, which imply progressively smaller sizes of the latent layers as they go deeper. Identifiability ensures consistent parameter estimation and inspires an interpretable design of the deep architecture. Computationally, we propose a scalable estimation pipeline of a layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm. This procedure can efficiently estimate models with exponentially many latent components. Extensive simulation studies validate our theoretical results and demonstrate the proposed algorithms’ excellent performance. We apply DDEs to three diverse real datasets for hierarchical topic modeling, image representation learning, response time modeling in educational testing, and obtain interpretable findings.
Corresponding Author: Yuqi Gu. Email: yuqi.gu@columbia.edu. Research is partially supported by NSF Grant DMS-2210796.
@@ 1 Introduction
The main contributions in this paper include rigorous identifiability theory and scalable computational pipelines for DDEs. For identifiability, we propose general identifiability conditions in terms of the dependence graph structure between adjacent layers in the graphical model (corresponding to the directed arrows in Figure 1). We work under the minimal possible assumptions in order to flexibly cover the various examples mentioned above. Next, we provide an informal statement of the identifiability conditions. Under an identifiable DDE, we also prove that a penalized-likelihood based estimator is consistent for estimating both the continuous parameters as well as the discrete graph structure.
@@
The DDE with a known number of latent variables is identifiable up to latent variable permutation in each layer, as long as each latent variable has at least three exclusive children. Under a weaker notion of “generic identifiability”, this condition can be relaxed to that each latent variable has at least three children that are not necessarily exclusive.
For computation, we propose a scalable estimation pipeline for DDEs. The multiple layers of nonlinear latent variables in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima. To address this challenging setting, our computational pipeline features a nuanced layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm. This procedure can efficiently handle models with a large number of latent variables. We achieve favorable simulation results for as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer, which amount to $2^{K^{(1)}}$ mixture components and define a very expressive model. Extensive simulation studies not only validate the identifiability results, but also demonstrate the excellent performance of the proposed algorithms. We apply DDEs to real data in three diverse tasks, including hierarchical topic modeling, image representation learning, and response time modeling in digital educational testing. Across these applications, DDEs extract highly interpretable latent structures and learn useful representations for downstream analyses.
We make brief remarks to place DDEs in the rapidly emerging field of generative AI. For the considered unsupervised learning setting, we use “generative model” to refer to probabilistic models where observed data are generated conditional on hidden variables. Classical examples with one hidden layer include mixture models and mixed membership models. Recently, powerful multilayer models have been proposed for more complex tasks, which either consist of multiple latent layers or a single latent layer transformed by deep neural networks. Popular models in machine learning include deep belief networks (Hinton et al.,, 2006), deep Boltzmann machines (Salakhutdinov and Hinton,, 2009), variational autoencoders (Ranzato and Szummer,, 2008), generative adversarial networks (Goodfellow et al.,, 2014), diffusion models (Sohl-Dickstein et al.,, 2015), and transformer-based models (Vaswani et al.,, 2017) such as large language models. See Section 2.2 for more discussions.
@@ 2 Deep Discrete Encoders
Notation. For a positive integer $K$, denote $[K]=\{1,\ldots,K\}$. For a matrix $\mathbf{G}$ with $J$ rows, let $\mathbf{g}_{1},\ldots,\mathbf{g}_{J}$ denote its row vectors. Let $\mathbf{0}_{K},\mathbf{1}_{K}$ be the all-zero vector and all-one vector of length $K$, respectively. Let $\mathbf{e}_{k}$ be the $K$-dimensional canonical basis vector. Let $\operatorname{\emph{g}_{\text{logistic}}}(x):={e^{x}}/({1+e^{x}})$ denote the logistic/sigmoid function. For a finite set $I$, let $S_{I}$ denote the collection of all permutation maps of $I$, and let $\text{id}_{I}$ be the identity permutation on $I$. For nonnegative sequences $\{(a_{N},b_{N})\}_{N\geq 1}$, write $a_{N}\ll b_{N}$ when ${a_{N}}/{b_{N}}\to 0$ as $N\to\infty$.
@@ 2.1 Model Definition
We formally define the $D$-latent-layer DDE as a generative model with $D$ discrete latent layers. For each $d\in[D]$, assume that the $(d-1)$th layer is generated conditional on the $d$th. Here, only the bottom layer (indexed by $d=0$) is observed, and all other layers are latent. The bottom layer data can take arbitrary values, but all latent variables are binary, similar to the celebrated deep belief networks (Hinton et al.,, 2006). Let $\mathbf{Y}=(Y_{1},\ldots,Y_{J})\in\times_{j=1}^{J}\mathcal{Y}_{j}$ denote the $J$-dimensional observed responses, where $\mathcal{Y}_{j}$ is the sample space for the $j$th response; see the end of this subsection for concrete examples. We work under the minimal assumption that each $\mathcal{Y}_{j}$ is a separable metric space. Denote the $d$th latent layer as $\mathbf{A}^{(d)}=\big{(}A_{1}^{(d)},\ldots,A_{K^{(d)}}^{(d)}\big{)}\in\{0,1\}^{K^{(d)}}$, which is a $K^{(d)}$-dimensional binary vector. A DDE has a shrinking-ladder-shaped architecture, with the dimension of each layer decreasing as $d$ increases: $K^{(D)}<\ldots<K^{(1)}<J$. See Figure 1 for a graphical model representation.
@@
Next, define the middle latent layers inductively as follows. For each $d>1$, suppose that $\mathbb{P}(\mathbf{A}^{(d)})$, the distribution of the $d$th layer, is given. Then, we define the $(d-1)$th layer distribution by assuming the conditional independence of $A_{1}^{(d-1)},\ldots,A_{K^{(d-1)}}^{(d-1)}$ given $\mathbf{A}^{(d)}$:
@@
Here, $\text{ParFam}_{j}$ denotes a specific parametric family, and let $H_{j}$ denote its parameter space. For convenience, let $p_{j}$ be the probability mass/density function of $\text{ParFam}_{j}$. The $g_{j}:\mathbb{R}\times[0,\infty)\rightarrow H_{j}$ is a known link function that maps the linear combinations to the parameters of the given parametric family. Here, $\gamma_{j}>0$ denotes the dispersion parameter, when there exists one. Throughout the paper, we will state all results under the more general assumption that there exists a dispersion parameter in the parametric family in (4). If not, one may ignore the notation $\gamma$. We further elaborate on the specific parameterizations in (4) for various response types $\mathcal{Y}_{j}$ at the end of this section.
@@
The $D$-latent-layer DDE is parametrized by $({\boldsymbol{p}},\mathcal{B},\boldsymbol{\gamma})$, where :=$\mathcal{B}:=\{\mathbf{B}^{(d)}\}_{d\in[D]}$ and $\boldsymbol{\gamma}:=(\gamma_{1},\ldots,\gamma_{J})$. Upon the above marginalization, the induced DDE is a highly expressive model with exponentially many latent mixture components. However, this expressivity also introduces identifiability and computation challenges, which we address in Sections 3 and 4.
In many scenarios, it is desirable for the coefficients $\mathcal{B}$ to be sparse, as this leads to a more interpretable and parsimonious data-generating mechanism. The interpretability stems from that if a latent variable is connected to only a few, rather than all, variables in the layer below, then these children variables can help pinpoint the interpretation of the latent parent. Similar sparse deep generative architectures have been considered in deep exponential families (Ranganath et al.,, 2015), Bayesian pyramids (Gu and Dunson,, 2023), and deep cognitive diagnostic models (Gu,, 2024). Additionally, as will be shown in Section 3.1, the sparsity of the coefficients play a key role in facilitating identifiability. To encode the sparsity of $\mathbf{B}^{(d)}$, for each $d\in[D]$, define a $K^{(d-1)}\times K^{(d)}$ binary matrix $\mathbf{G}^{(d)}=(g_{k,l})$, where $g_{k,l}=1$ if the corresponding coefficient $\beta_{k,l}^{(d)}$ is nonzero, and 0 otherwise. Define $K^{(0)}:=J$ for notational convenience. By (2) and (3), $\mathbf{G}^{(d)}$ can also be viewed as the adjacency matrix or “graphical matrix” between the $(d-1)$th layer and the $d$th layer in the graphical model representation (see Figure 1). Collect all graphical matrices $\mathbf{G}^{(d)}$s by defining $\mathcal{G}:=\{\mathbf{G}^{(d)}\}_{d\in[D]}$. Now, we formally define $D$-latent-layer DDEs, which also incorporate $\mathcal{G}$ as unknown parameters.
@@ 2.2 Connections with Existing Models
Deep Probabilistic Graphical Models in Machine Learning.
There are many deep latent variable models in the machine learning literature, including the popular deep Boltzmann machine (DBM, Salakhutdinov and Hinton,, 2009), deep belief networks (DBNs, Hinton et al.,, 2006), and deep exponential families (DEFs, Ranganath et al.,, 2015). The DBM and DBN both contain multiple binary latent layers and differ in the directions of the edges; see Figure 2. DBM and DBN have been originally proposed to model binary data and have been later extended to handle continuous or count responses (Cho et al.,, 2013; Gan et al.,, 2015; Li et al.,, 2019). We recommend the review Salakhutdinov, (2015) for more details and references on DBM and DBNs. DEFs are an unsupervised modeling framework that uses exponential families to model conditional distributions for each layer.
[…]
@@
As to be shown later, DDEs resolve all these issues by assuming a shrinking-ladder architecture of entirely discrete latent variables and potentially sparse layerwise connections (see Figure 2). This allows us to establish identifiability and consistency as well as effectively reduce the model dimension and interpret the latent structure. Recently, Gu and Dunson, (2023) proposed Bayesian pyramids, which are identifiable multilayer discrete latent variable models with a pyramid structure. But the methodology and identifiability theory therein are restricted to multivariate categorical data. DDEs significantly broaden the applicability of Bayesian pyramids by modeling arbitrarily flexible data distributions while still remaining identifiable. For computation, we propose scalable frequentist estimation approaches for DDEs, rather than adopting Bayesian MCMC inference as in Gu and Dunson, (2023).
Psychometrics.
3 Theoretical Guarantees for General DDEs
@@ 3.1 Model Identifiability
In this section, we establish the identifiability of DDEs. Assuming known numbers of latent variables, we prove that the continuous model parameters and the discrete graph structures between layers can be uniquely identified under certain conditions on the true graphical matrices $\mathbf{G}^{(d)}$’s. We first make an assumption to address some trivial ambiguities.
@@
Assume that the graphical matrices $\mathcal{G}=\{\mathbf{G}^{(d)}\}_{d\in[D]}$ and proportion parameters $\boldsymbol{p}$ satisfy the following conditions. Here, for notational convenience, let $K^{(0)}:=J$.
@@
For a D-layer DDE with $\mathcal{K}$ latent variables, we define an equivalence relationship “$\sim_{\mathcal{K}}$” by setting $(\boldsymbol{\Theta},\mathcal{G})\sim_{\mathcal{K}}(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathcal{G}})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exists permutations $\sigma^{(d)}\in S_{[K^{(d)}]}$ for all $d\in[D]$ such that the following hold:
@@
Despite the seemingly heavy notation, the equivalence relationship is quite natural. Here, each permutation $\sigma^{(d)}$ addresses the label-switching issue for indexing the latent variables in the $d$th latent layer. Below is a main result on the identifiability of $D$-latent-layer DDEs.
@@
Consider the $D$-latent-layer DDE with $\mathcal{K}$ latent variables and true parameters $(\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star})$. Suppose that for all $d\leq D-1$, the true graphical matrices $\mathbf{G}^{(d)\star}$ and parameters $\mathbf{B}^{(d)\star}$ satisfy the following conditions:
@@
    $\mathbf{G}^{(d)\star}$ contains two identity matrices after permuting the rows. Without the loss of generality, suppose that the first $2K^{(d)}$ rows of $\mathbf{G}^{(d)\star}$ are $\big{[}\mathbf{I}_{K^{(d)}},\mathbf{I}_{K^{(d)}}\big{]}^{\top}.$
@@
    For any $\boldsymbol{\alpha}^{(d)}\neq\boldsymbol{\alpha}^{\prime(d)}\in\{0,1\}^{K^{(d)}}$, there exists $j>2K^{(d)}$ such that $\sum_{k=1}^{K^{(d)}}\beta_{j,k}^{(d)\star}(\alpha_{k}^{(d)}-\alpha_{k}^{(d)^{\prime}})\neq 0$.
@@
Our key observation behind proving the identifiability of the complex deep generative structures in DDEs is that, with discrete latent layers, it suffices to establish identifiability for a one-latent-layer model and proceed in a layer-wise manner inductively. To elaborate, consider a “collapsed” DDE where the latent layers indexed by $d=2,\ldots,D$ are marginalized out to give a probability mass function for the first latent layer: $\{\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}):\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}\}$. If the collapsed model is proven to be identifiable, that means the conditional distributions $\mathbb{P}(Y_{j}\mid\mathbf{A}^{(1)})$ and the marginal distribution $\mathbb{P}(\mathbf{A}^{(1)})$ can be uniquely identified from the data distribution $\mathbb{P}(\mathbf{Y})$. In this case, we can determine all values of $\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})$ for every $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$. Then, we can theoretically treat the shallowest latent layer $\mathbf{A}^{(1)}$ as if it were observed, because its probability mass function (pmf) is now identified and known. Then, viewing this pmf as the marginal distribution of $K^{(1)}$-dimensional observed variables of a $(D-1)$-latent-layer Bernoulli-DDE, we can inductively identify all parameters via a layerwise argument. Please see Supplementary Material S.1 for the detailed proof.
We interpret the conditions in Theorem 1. Condition A requires each latent variable to have at least two exclusive children in the layer below. Condition B is more technical and is introduced to distinguish the different binary latent configurations $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}$. Condition B holds when $\mathbf{G}^{(d)\star}$ also contains exclusive children for all latent variables.
The exclusive children requirements in condition A can be further relaxed under a weaker notion of generic identifiability. Generic identifiability allows a measure-zero subset of the parameter space to be non-identifiable, and often holds under weaker conditions than that in Definition 3. As the concept was originally proposed under a continuous parameter space (Allman et al.,, 2009), we consider the following modified definition that considers a smaller parameter space for the coefficients $\mathcal{B}$ given the true graphical matrices.
@@
Consider a $D$-latent-layer DDE with $\mathcal{K}$ latent variables, graphical matrices $\mathcal{G}^{\star}$, and true parameters belonging to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$. The model is generically identifiable up to $\sim_{\mathcal{K}}$ when
[…]
is a measure-zero set with respect to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
We introduce an additional assumption on the parametric families used to model the conditional distribution $Y_{j}\mid\mathbf{A}^{(1)}$ in (4). This is a technical assumption that arises from our proof technique for dealing with measure-zero sets. This assumption holds for all example parametric families described in Section 2.1, and more generally for exponential families with an analytic log-partition function.
@@
Now, we are ready to state the generic identifiability result for $D$-latent-layer DDEs.
@@
Consider the $D$-latent-layer DDE with $\mathcal{K}$ latent variables and true graphical matrices $\mathcal{G}^{\star}$, where all parametric families and link function $g_{j}$s are analytic, and the true parameter lives in $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$. Suppose that for all $d\leq D-1$, the true graphical matrices $\mathbf{G}^{(d)}$ satisfy the following condition C:

### v2

@@ Abstract
In the era of generative AI, deep generative models (DGMs) with latent representations have gained tremendous popularity. Despite their impressive empirical performance, the statistical properties of these models remain underexplored. DGMs are often overparametrized, non-identifiable, and uninterpretable black boxes, raising serious concerns when deploying them in high-stakes applications. Motivated by this, we propose interpretable deep generative models for rich data types with discrete latent layers, called Deep Discrete Encoders (DDEs). A DDE is a directed graphical model with multiple binary latent layers. Theoretically, we propose transparent identifiability conditions for DDEs, which imply progressively smaller sizes of the latent layers as they go deeper. Identifiability ensures consistent parameter estimation and inspires an interpretable design of the deep architecture. Computationally, we propose a scalable estimation pipeline of a layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm. This procedure can efficiently estimate models with exponentially many latent components. Extensive simulation studies for high-dimensional data and deep architectures validate our theoretical results and demonstrate the excellent performance of our algorithms. We apply DDEs to three diverse real datasets with different data types to perform hierarchical topic modeling, image representation learning, and response time modeling in educational testing.
Corresponding Author: Yuqi Gu. Email: yuqi.gu@columbia.edu.
@@ 1 Introduction
The main contributions of this paper include rigorous identifiability theory and scalable computational pipelines for DDEs. For identifiability, we propose general identifiability conditions in terms of the probabilistic graph structures between layers in the graphical model (corresponding to the directed arrows in Figure 1). We work under the minimal possible assumptions in order to flexibly cover the various examples and data types mentioned above. Next, we provide an informal statement of the identifiability conditions. Under an identifiable DDE, we also prove that a penalized-likelihood based estimator is consistent for estimating both the continuous parameters as well as the discrete graph structure.
@@
The DDE is identifiable up to latent variable permutation in each layer, as long as each latent variable has at least three pure children (which only has one parent variable). Under a weaker notion of “generic identifiability”, this condition can be relaxed to that each latent variable has at least three children that are not necessarily pure.
For computation, we propose a scalable estimation pipeline for DDEs. The multiple layers of nonlinear latent variables in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima. To address this challenging issue, our computational pipeline features a nuanced layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm. This procedure can efficiently handle models with a large number of latent variables. We achieve favorable simulation results for as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer, which amount to $2^{K^{(1)}}$ mixture components and define a very expressive model. Extensive simulation studies not only validate the identifiability results, but also demonstrate the excellent performance of the proposed algorithms. We apply DDEs to real data in three diverse tasks, including hierarchical topic modeling, image representation learning, and multimodal modeling in digital educational testing. Across these applications, DDEs extract highly interpretable latent structures and learn useful representations for downstream analyses.
We make brief remarks to place DDEs in the rapidly emerging field of generative AI. For the considered unsupervised learning setting, we use “generative model” to refer to probabilistic models where observed data are generated conditional on hidden variables. Recently, powerful multilayer models have been proposed for more complex tasks, which either consist of multiple latent layers or a single latent layer transformed by deep neural networks. Popular models in machine learning include deep belief networks (DBNs, Hinton et al., , 2006), deep Boltzmann machines (DBMs, Salakhutdinov and Hinton, , 2009), variational autoencoders (Ranzato and Szummer, , 2008), generative adversarial networks (Goodfellow et al., , 2014), diffusion models (Sohl-Dickstein et al., , 2015), and transformer-based models (Vaswani et al., , 2017) such as large language models. Our proposed models are most closely related to DBNs and DBMs. See Section 2.2 for more discussions.
@@ 2 Deep Discrete Encoders
Notation. For a positive integer $K$, denote $K]=\{1,\ldots,K\}$. For a matrix $\mathbf{G}$ with $J$ rows, let $\mathbf{g}_{1},\ldots,\mathbf{g}_{J}$ denote its row vectors. Let $\mathbf{0}_{K},\mathbf{1}_{K}$ be the all-zero vector and all-one vector of length $K$, respectively. Let $\mathbf{e}_{k}$ be the $K$-dimensional canonical basis vector. Let $\operatorname{\emph{g}_{\text{logistic}}}(x):={e^{x}}/({1+e^{x}})$ denote the logistic/sigmoid function. For a finite set $I$, let $S_{I}$ denote the collection of all permutation maps of $I$, and let $\text{id}_{I}$ be the identity permutation on $I$.
@@ 2.1 Model Definition
The $D$-latent-layer DDE is a generative model with $D$ discrete latent layers. For each $d\in[D]$, assume that the (d1)(d-1)th layer is generated conditional on the $d$th. Here, only the bottom layer (indexed by $d=0$) is observed, and all other layers are latent. The bottom layer data can take arbitrary values, but all latent variables are binary, similar to the celebrated deep belief networks (Hinton et al., , 2006). Let $\mathbf{Y}=(Y_{1},\ldots,Y_{J})\in\times_{j=1}^{J}\mathcal{Y}_{j}$ denote the $J$-dimensional observed responses, where $\mathcal{Y}_{j}$ is the sample space for the $j$th response; see the end of this subsection for concrete examples. We work under the minimal assumption that each $\mathcal{Y}_{j}$ is a separable metric space. Denote the $d$th latent layer as $\mathbf{A}^{(d)}=\big{(}A_{1}^{(d)},\ldots,A_{K^{(d)}}^{(d)}\big{)}\in\{0,1\}^{K^{(d)}}$, which is a $K^{(d)}$-dimensional binary vector. A DDE has a shrinking-ladder-shaped deep architecture, with the dimension of each layer decreasing as $d$ increases: $K^{(D)}<\ldots<K^{(1)}<J$. See Figure 1 for a graphical model representation.
@@
Next, define the middle latent layers inductively as follows. For each $d>1$, suppose that $\mathbb{P}(\mathbf{A}^{(d)})$, the distribution of the $d$th layer, is given. Then, we define the (d1)(d-1)th layer distribution by assuming the conditional independence of $A_{1}^{(d-1)},\ldots,A_{K^{(d-1)}}^{(d-1)}$ given $\mathbf{A}^{(d)}$:
@@
Here, $\text{ParFam}_{j}$ denotes a specific identifiable parametric family, and let $H_{j}$ denote its parameter space. For convenience, let $p_{j}$ be the probability mass/density function of $\text{ParFam}_{j}$. The $g_{j}:\mathbb{R}\times[0,\infty)\rightarrow H_{j}$ is a known link function that maps the linear combinations to the parameters of the given parametric family. Here, $\gamma_{j}>0$ denotes the dispersion parameter, when there exists one. Throughout the paper, we will state all results under the more general assumption that there exists a dispersion parameter in the parametric family in (4). If not, one may ignore the notation $\gamma$. We further elaborate on the specific parameterizations in (4) for various response types $\mathcal{Y}_{j}$ at the end of this section.
@@
The $D$-latent-layer DDE is parametrized by $({\boldsymbol{p}},\mathcal{B},\boldsymbol{\gamma})$, where $\mathcal{B}:=\{\mathbf{B}^{(d)}\}_{d\in[D]}$ and $\boldsymbol{\gamma}:=(\gamma_{1},\ldots,\gamma_{J})$. Upon the above marginalization, the induced DDE is a highly expressive model with exponentially many latent mixture components. However, this expressivity also introduces identifiability and computation challenges, which we address in Sections 3 and 4.
In many scenarios, it is desirable for the coefficients $\mathcal{B}$ to be sparse, as this leads to a more interpretable and parsimonious data-generating mechanism. The interpretability stems from that if a latent variable is connected to only a few, rather than all, variables in the layer below, then these children variables can help pinpoint the interpretation of the latent parent. Similar sparse deep generative architectures have been considered in deep exponential families (Ranganath et al., , 2015), Bayesian pyramids (Gu and Dunson, , 2023), and deep cognitive diagnostic models (Gu, , 2024). Additionally, as will be shown in Section 3.1, the sparsity of the coefficients play a key role in facilitating identifiability. To encode the sparsity of $\mathbf{B}^{(d)}$, for each $d\in[D]$, define a $K^{(d-1)}\times K^{(d)}$ binary matrix $\mathbf{G}^{(d)}=(g_{k,l})$, where $g_{k,l}=1$ if the corresponding coefficient $\beta_{k,l}^{(d)}$ is nonzero, and 0 otherwise. Define $K^{(0)}:=J$ for notational convenience. By (2) and (3), $\mathbf{G}^{(d)}$ can also be viewed as the adjacency matrix or “graphical matrix” between the (d1)(d-1)th layer and the $d$th layer in the graphical model representation (see Figure 1). Collect all $\mathbf{G}^{(d)}$s by defining $\mathcal{G}:=\{\mathbf{G}^{(d)}\}_{d\in[D]}$, and collect the number of latent variables in each layer by defining $\mathcal{K}:=\{K^{(d)}\}_{d\in[D]}.$ Now, we formally define $D$-latent-layer DDEs, which also incorporate $\mathcal{G}$ as unknown parameters.
@@ 2.2 Connections with Existing Models
Related deep latent variable models in the machine learning literature include the deep Boltzmann machine (DBM, Salakhutdinov and Hinton, , 2009), deep belief networks (DBNs, Hinton et al., , 2006), and deep exponential families (DEFs, Ranganath et al., , 2015). The DBM and DBN both contain multiple binary latent layers and differ in the directions of the edges; see Figure 2. DBM and DBN have been originally proposed to model binary data and have been later extended to handle continuous or count responses (Cho et al., , 2013; Gan et al., , 2015; Li et al., , 2019). We recommend the review Salakhutdinov, (2015) for more details and references on DBM and DBNs. DEFs are an unsupervised modeling framework that uses exponential families to model conditional distributions for each layer.
$A_{1}^{(2)}$$A_{2}^{(2)}$$A_{3}^{(2)}$$A_{1}^{(1)}$$A_{2}^{(1)}$$A_{3}^{(1)}$$Y_{1}$$Y_{2}$$Y_{3}$
@@
As to be shown later, DDEs resolve all these issues by assuming a shrinking-ladder architecture of entirely discrete latent variables and potentially sparse layerwise connections (see Figure 2). This allows us to establish identifiability and consistency as well as effectively reduce the model dimension and interpret the latent structure.
On a related note, several identifiable deep models have recently been proposed. Gu and Dunson, (2023) proposed Bayesian pyramids, which are identifiable multilayer discrete latent variable models with a pyramid structure. But the methodology and identifiability theory therein are restricted to multivariate categorical data. DDEs significantly broaden the applicability of Bayesian pyramids by modeling arbitrarily flexible data types while still remaining identifiable. Kong et al., (2024) considered a modeling framework allowing both discrete and continuous latent variables, where the discrete part of the model can have more flexible graphical structures than the multi-layer structure. However, as a tradeoff, they consider a weaker notion of identifiability (up to atomic cover structures) and propose less explicit conditions for identifiability. Finally, Anandkumar et al., (2013) and Xie et al., (2024) consider continuous latent variable models and prove identifiability under potentially more general graph structures, but assume that all conditional distributions are linear. In contrast, DDEs allow arbitrary nonlinear link functions, which improves representation power. We remark that DDEs are identifiable even when higher-order interaction effects among latent variables are in the model (as detailed in Supplement S.1.6). We provide additional literature review regarding identifiable VAEs and psychometric models in Supplement S.6.
3 Theoretical Guarantees for DDEs
@@ 3.1 Model Identifiability
In this section, we establish the identifiability of DDEs. Assuming known numbers of latent variables, we prove that both the continuous model parameters and the discrete graph structures between layers can be uniquely identified under certain conditions on the true graphical matrices $\mathbf{G}^{(d)}$’s. We first make an assumption to address some trivial ambiguities. For notational convenience, denote $K^{(0)}:=J$ and $A_{j}^{(0)}:=Y_{j}$.
@@
Assume that the graphical matrices $\mathcal{G}=\{\mathbf{G}^{(d)}\}_{d\in[D]}$ and proportion parameters $\boldsymbol{p}$ satisfy the following conditions.
@@
[…]
@@
Despite the seemingly heavy notation, the equivalence relationship is quite natural. For example, in the right panel of Figure 1, the location of the latent variables in each layer can be arbitrarily permuted without changing the likelihood. In other words, the latent variable “sports” and “car” can be equivalently indexed by $(1,2)$ or $(2,1)$ as long as the associated arrows are permuted accordingly. For the $d$th latent layer, the permutation $\sigma^{(d)}$ corresponds to this fundamental yet relatively trivial label-switching map. The following theorem is our first main result on the identifiability of DDEs. Here, we say that a variable $X$ is a “pure child” of $Y$ if $Y$ is the only parent of $X$.
@@
Consider a $D$-latent-layer DDE with true parameters $(\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star})$, where the number of latent variables, $\mathcal{K}$, is given. Suppose that for all $d\leq D-1$, the true graphical structures $\mathbf{G}^{(d)\star}$ and parameters $\mathbf{B}^{(d)\star}$ satisfy the following conditions:
@@
    Each latent variable $A_{k}^{(d)}$ has at least two pure children $A_{j_{k,1}}^{(d-1)},A_{j_{k,2}}^{(d-1)}$ according to $\mathbf{G}^{(d)\star}$.
@@
    For any $\boldsymbol{\alpha}^{(d)}\neq\boldsymbol{\alpha}^{\prime(d)}\in\{0,1\}^{K^{(d)}}$, there exists $j\in[K^{(d-1)}]\setminus\cup_{k=1}^{K^{(d)}}\{j_{k,1},j_{k,2}\}$ such that $\sum_{k=1}^{K^{(d)}}\beta_{j,k}^{(d)\star}(\alpha_{k}^{(d)}-\alpha_{k}^{{}^{\prime}(d)})\neq 0.$
@@
Our key observation behind proving the identifiability of the complex deep generative structures in DDEs is that, with discrete latent layers, it suffices to establish identifiability for a one-latent-layer model and proceed in a layer-wise manner inductively. To elaborate, consider a “collapsed” DDE where the latent layers indexed by $d=2,\ldots,D$ are marginalized out to give a probability mass function (pmf) for the first latent layer: $\{\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}):\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}\}$. If the collapsed model is proven to be identifiable, that means the conditional distributions $\mathbb{P}(Y_{j}\mid\mathbf{A}^{(1)})$ and the marginal distribution $\mathbb{P}(\mathbf{A}^{(1)})$ can be uniquely identified from the data distribution $\mathbb{P}(\mathbf{Y})$. In this case, we can determine the structured pmf up to the inevitable permutation of the latent variables (indexed by $\sigma^{(1)}\in S_{[K^{(1)}]}$ in Definition 3). Then, we can theoretically treat the shallowest latent layer $\mathbf{A}^{(1)}$ as if it was observed, because its probability mass function is now identified and known. Then, viewing this pmf as the marginal distribution of $K^{(1)}$-dimensional observed variables of a (D1)(D-1)-latent-layer Bernoulli-DDE, we can inductively identify all parameters via a layerwise argument. Please see Supplementary Material S.1 for the detailed proof.
We next give an interpretation of the conditions in Theorem 1. Condition A requires each latent variable $A_{k}^{(d)}$ to have at least two pure children in the layer below. For example, the latent variable “tech” in the right panel of Figure 1 has two pure children “graphics” and “space”. Condition B is more technical and is introduced to distinguish the different binary latent configurations $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}$. For example, condition B holds when each latent variable $A_{k}^{(d)}$ has a third pure child that is distinct from those in condition A. We also provide identifiability guarantees for the number of latent variables in Supplement S.1.6.
The pure children requirements in condition A can be further relaxed under a weaker notion of generic identifiability. Generic identifiability allows a measure-zero subset of the parameter space to be non-identifiable, and often holds under weaker conditions than that in Definition 3. As the concept was originally proposed under a continuous parameter space (Allman et al., , 2009), we consider the following modified definition that considers a smaller parameter space for the coefficients $\mathcal{B}$ given the true graphical matrices.
@@
Consider a $D$-latent-layer DDE with $\mathcal{K}$ latent variables, graphical matrices $\mathcal{G}^{\star}$, and true parameters belonging to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$. The model is generically identifiable up to $\sim_{\mathcal{K}}$ when $\{\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$: there exists $(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathcal{G}})\not\sim_{\mathcal{K}}(\boldsymbol{\Theta},\mathcal{G}^{\star})$ such that $\mathbb{P}_{\widetilde{\boldsymbol{\Theta}},\widetilde{\mathcal{G}}}=\mathbb{P}_{\boldsymbol{\Theta},\mathcal{G}^{\star}}\}$ is a measure-zero set with respect to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
For generic identifiability, we introduce an additional assumption on the parametric families used to model the conditional distribution $Y_{j}\mid\mathbf{A}^{(1)}$ in (4). This is a technical assumption that arises from our proof technique for dealing with measure-zero sets. This assumption holds for all example parametric families described in Section 2.1, and more generally for exponential families with an analytic log-partition function.
@@
We next state the generic identifiability result for DDEs. In graph theory, a bipartite graph is said to have a “perfect matching” if it contains a set of edges without common vertices that covers every vertex of the graph (Hall, , 2011); see Example 1 for an illustration.
@@
Consider the $D$-latent-layer DDE where the number of latent variables $\mathcal{K}$ is given, and all parametric families and link function $g_{j}$s are analytic. Let $\mathcal{G}^{\star}$ denote the true graphical matrices and suppose the true parameter lives in $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$. Suppose that for all $d\leq D-1$, the true graphical structure $\mathbf{G}^{(d)}$ satisfy the following condition C: